A first order reaction has a half life period of 69.3 sec. At 0.10 mol L^-1 reactant concentration, the rate will be-
a.10^-4 M sec^-1     b.10^-3 M sec^-1

Grade:

1 Answers

Arun
5 years ago
Rate Constant, k = 0.693/t0.5 = 0.693/69.3 = 0.01
 
Rate = k[A]1 = 0.01 × 0.1 = 10-3 mol/l/sec
 
 
Regards
Arun (askIITians forum expert)